The sum of two numbers is -22. The difference of the two numbers is 8. What are the two numbers?
Answer by ikleyn(53748)

.

    x + y = -22

    x - y =   8
 ------------------- Add these equations to get

    2x = -22 + 8 = -14  --->  x = -14/2 = -7;  y = -22 - x = -22 - (-7) = -22 + 7 = -15.


ANSWER.  The numbers are -7 and -15.

Solved.

There are other ways to solve, but this one is the most educative (in my view).